### v1

# Ergodic Estimates of One-Step Numerical Approximations for Superlinear SODEs

Abstract.
This paper establishes the first-order convergence rate for the ergodic error of numerical approximations to a class of stochastic ODEs (SODEs) with superlinear coefficients and multiplicative noise. By leveraging the generator approach to the Stein method, we derive a general error representation formula for one-step numerical schemes. Under suitable dissipativity and smoothness conditions, we prove that the error between the accurate invariant measure $\pi$ and the numerical invariant measure $\pi_{\tau}$ is of order $\mathscr{O}(\tau)$, which is sharp. Our framework applies to several recently studied schemes, including the tamed Euler, projected Euler, and backward Euler methods.
Key words and phrases:
Ergodic estimate, Stein method, stochastic one-step approximations, superlinear SODEs
2010 Mathematics Subject Classification:
60H35; 37M25; 65C30
The authors are supported by the National Natural Science Foundation of China, No. 12101296, Guangdong Basic and Applied Basic Research Foundation, No. 2024A1515012348, and Shenzhen Basic Research Special Project (Natural Science Foundation) Basic Research (General Project), Nos. JCYJ20220530112814033 and JCYJ20240813094919026.
1. Introduction
Consider the following $d$-dimensional SODE driven by an $m$-dimensional Wiener process $W$ in a complete filtered probability space $(\Omega,\mathscr{F},\{\mathscr{F}_{t}\}_{t\geq 0},\mathbb{P})$:
(1.1) $\mathrm{d}X_{t}=b(X_{t})\,\mathrm{d}t+\sigma(X_{t})\,\mathrm{d}W(t),\quad t>0.$
Here $b:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ and $\sigma:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times m}$ are measurable functions satisfying certain monotone and polynomial growth conditions.
Under appropriate conditions, Eq. (1.1) admits a unique invariant measure $\pi$. While characterizing $\pi$ explicitly is generally infeasible, it is known only for specific cases, such as the gradient Langevin equation with additive noise or one-dimensional SODEs with multiplicative nondegenerate noise [5]. Therefore, constructing accurate numerical approximations of $\pi$ is a problem of significant interest. The study of long-time approximations for SODEs is a growing interest, with significant applications in scientific computing, statistics, machine learning, and generative AI. For instance, sampling in score-based diffusion models–a key technique in generative AI–fundamentally relies on long-time SODE approximations.
The Euler–Maruyama (EM) scheme is effective for Lipschitz Eq. (1.1) (see, e.g., [2, 16]) or Markovian switching diffusion (see, e.g., [17]). However, for coefficients with superlinear growth, the EM scheme would blow up in $p$-th moment for all $p\geq 2$ [3], precluding ergodicity. Consequently, several modified schemes have been developed to preserve ergodicity, including the backward Euler method (BEM) and stochastic theta method [7, 10, 14], and the modified Euler method (MEM)–including tamed Euler method (TEM) and projected Euler method (PEM) [12, 13]; see also [6] for Markovian switching diffusions.
A central question is to quantify the ergodic error $|\pi(\varphi)-\pi_{\tau}(\varphi)|$ between $\pi_{\tau}$ and $\pi$ for a given step-size $\tau\in(0,1)$ and a test function $\varphi$. Using the generator approach [1] to the Stein method [18], the authors of [2] derived an ergodic error representation for additive noise cases and obtained a rate of $\mathscr{O}(\tau^{1/2}|\ln\tau|)$ for the EM scheme (the so-called unadjusted Langevin algorithm in the Langenvin equation case) with Lipschitz drift and nondegenerate additive noise, where $\varphi$ is a Lipschitz function. This was later generalized in [4] to dissipative Lipschitz systems with nondegenerate multiplicative noise and jump diffusions, removing the logarithmic factor for test functions in $\mathcal{C}_{b}^{3}(\mathbb{R}^{d})$. We also refer to [15] and references therein for nonasymptotic bound analysis for the decreasing-step EM scheme.
Since the weak convergence rate is typically twice the strong convergence rate, an ergodic error of order $\mathscr{O}(\tau)$ is expected. This paper achieves this optimal first-order rate for a family of superlinear SODEs with multiplicative noise and a class of one-step numerical approximations, assuming test functions in $\mathcal{C}_{b}^{4}(\mathbb{R}^{d})$.
Our contributions:
  1. (1)
    We derive the Stein equation with an explicit solution linked to the invariant measure under fewer conditions and introduce a new representation of the ergodic error for general one-step numerical approximations, decomposing it into an initial error and a remainder from the continuous interpolation (see Theorems 2.1 and 2.3, respectively).
  2. (2)
    Our framework applies to several recently studied numerical schemes–including the tamed Euler, projected Euler, and backward Euler methods–and establishes the optimal first order of convergence (see Theorem 3.5).
The paper is organized as follows. Section 2 presents the main ideas, including the Stein equation and the general ergodic error estimate. Section 3 applies these results to superlinear SODEs and provides detailed analyses for specific numerical schemes.
2. Main Idea and Results
This section will introduce some frequently used notations, the Stein equation with its solution, and the main result on the ergodic error for general one-step approximations.
2.1. Notations
Let us begin with some frequently used notations. Throughout this paper, we denote by $\left\lvert\cdot\right\rvert$, $\langle\cdot,\cdot\rangle$, and $\left\lVert\cdot\right\rVert_{HS}$ the Euclidean norm, the inner product in $\mathbb{R}^{d}$, and the Hilbert–Schmidt norm in $\mathbb{R}^{d\times m}$.
In addition, we use $\mathcal{B}_{b}(\mathbb{R}^{d})$ and ${\rm Lip}(\mathbb{R}^{d})$ to denote the Banach spaces of all bounded measurable and Lipschitz continuous mappings $\phi:\mathbb{R}^{d}\rightarrow\mathbb{R}$. Moreover, for $k\in\mathbb{N}_{+}$, denote by $C_{b}^{k}(\mathbb{R}^{d})$ the Banach spaces of all continuous functions with bounded partial derivatives $\nabla^{i}\varphi(x)\in\mathcal{L}((\mathbb{R}^{d})^{\otimes i};\mathbb{R})$, the $i$-fold linear functional in $\mathbb{R}^{d}$ with $1\leq i\leq k$, endowed with the seminorm
$|\varphi|_{k}:=\sum_{i=1}^{k}\sup_{x\in\mathbb{R}^{d}}\|\nabla^{i}\varphi(x)\|_{\mathcal{L}((\mathbb{R}^{d})^{\otimes i};\mathbb{R})}.$
2.2. Stein Equation
Assume Eq. (1.1) has a unique continuous strong solution $(X_{t}^{\cdot})_{t\geq 0}$ with Markov semigroup $(P_{t})_{t\geq 0}$ given by
(2.1) $P_{t}\varphi(x)=\mathbb{E}\varphi(X_{t}^{x}),\quad x\in\mathbb{R},~\varphi\in\mathcal{B}_{b}(\mathbb{R}^{d}),~t\geq 0.$
Assume $(X_{t}^{\cdot})_{t\geq 0}$ has bounded algebraic moments in the sense
$\sup_{t\geq 0}\mathbb{E}|X_{t}^{x}|^{p}\leq C(1+|x|^{p}),\quad p\geq 2,~x\in\mathbb{R}^{d},$
and $\mathbb{E}\int_{0}^{T}|\sigma_{j}(X_{t}^{\cdot})|^{2}\mathrm{d}t<\infty$ for all $T>0$ and $j=\{1,2,\cdots,m\}$, so that $\int_{0}^{\cdot}\sigma(X_{s}^{\cdot})\mathrm{d}W(s)$ is a martingale. We note that the definition (2.1) makes sense for $\varphi\in{\rm Lip}(\mathbb{R}^{d})$. Assume $(P_{t})_{t\geq 0}$ has a unique invariant measure $\pi$ with bounded algebraic moments and mixing property:
(2.2) $\lim_{t\to+\infty}P_{t}\varphi(x)=\pi(\varphi)=:\int_{\mathbb{R}^{d}}\varphi(x)\pi(\mathrm{d}x),\quad x\in\mathbb{R},~\varphi\in\mathcal{B}_{b}(\mathbb{R}^{d}).$
For an $f\in C^{2}(\mathbb{R}^{d})$, Itô formula gives
(2.3) $\mathrm{d}f(X_{t})$ $=\mathcal{A}f(X_{t})\mathrm{d}t+\nabla f(X_{t})\sigma(X_{t})\mathrm{d}W(t),\quad t>0,$
where $\mathcal{A}$ is the infinitesimal generator of Eq. (1.1):
(2.4) $\mathcal{A}f(x):=\nabla f(x)b(x)+\frac{1}{2}\sum_{j=1}^{m}\nabla^{2}f(x)(\sigma_{j}(x),\sigma_{j}(x)),\quad x\in\mathbb{R}^{d}.$
Then (2.3) yields that
(2.5) $\mathbb{E}[f(X_{t})-f(X_{0})]=\int_{0}^{t}\mathbb{E}\mathcal{A}f(X_{s})\mathrm{d}s,\quad t\geq 0.$
If $X_{0}\sim\pi$, we have $\int_{0}^{t}\mathbb{E}[\mathcal{A}f(X_{s})]\mathrm{d}s=0$ for any $t\geq 0$. Thus $\mathbb{E}[\mathcal{A}f(X_{0})]=0$ by the continuity of the process $X$.
Given an $\varphi\in\mathcal{B}_{b}(\mathbb{R}^{d})$ or ${\rm Lip}(\mathbb{R}^{d})$, let us consider the Stein equation associated with the invariant measure $\pi$ of Eq. (1.1):
(2.6) $\varphi(x)-\pi(\varphi)=\mathcal{A}f_{\varphi}(x),\quad x\in\mathbb{R}^{d}.$
The following theorem constructs a solution to the Stein equation (2.6).
Theorem 2.1.
For any $\varphi\in\mathcal{B}_{b}(\mathbb{R}^{d})$ or ${\rm Lip}(\mathbb{R}^{d})$ such that $P_{\cdot}\varphi\in L^{1}(0,\infty;\mathcal{C}_{b}^{2}(\mathbb{R}^{d}))$, Eq. (2.6) admits a solution $f_{\varphi}\in\mathcal{C}_{b}^{2}(\mathbb{R}^{d})$ given by
(2.7) $f_{\varphi}=-\int_{0}^{+\infty}[P_{t}\varphi-\pi(\varphi)]\mathrm{d}t.$
Proof.
Let $T>0$. Applying $P_{t}$ to both sides of Eq. (2.6), integrating over $t\in[0,T]$, and using the definition (2.1) and the identity (2.5), we obtain
(2.8) $\begin{aligned} \int_{0}^{T}P_{t}[\varphi(x)-\pi(\varphi)]\mathrm{d}t&=\int_{0}^{T}P_{t}\mathcal{A}f_{\varphi}(x)\mathrm{d}t=\int_{0}^{T}\mathbb{E}[\mathcal{A}f_{\varphi}(X_{t}^{x})]\mathrm{d}t\\ &=\mathbb{E}[f_{\varphi}(X_{T}^{x})-f_{\varphi}(X_{0}^{x})]=P_{T}f_{\varphi}(x)-f_{\varphi}(x).\end{aligned}$
Letting $T\to+\infty$ and using the mixing property (2.2), we derive
$f_{\varphi}(x)=-\int_{0}^{+\infty}P_{t}[\varphi(x)-\pi(\varphi)]\mathrm{d}t+\pi(f_{\varphi}),$
so that $f_{\varphi}$ given in (2.7) is a solution of the Stein equation (2.6). ∎
Remark 2.2.
The solution of the Stein equation (2.6) is not unique up to a constant. Indeed, in the case $d=m=1$, any function $f:=f_{\varphi}+f_{b,\sigma}$ with $f_{b,\sigma}^{\prime}(x)=c_{0}\exp(\int_{0}^{x}-\frac{2b(y)}{\sigma(y)^{2}}\mathrm{d}y)$ for any constant $c_{0}$ is also a solution of Eq. (2.6).
In the rest of this paper, we always take $f_{\varphi}$ as in (2.7).
2.3. Ergodic Estimate of One-step Approximations
Let $\tau\in(0,1)$ be a fixed step-size, $t_{k}:=\tau k$, and $\delta_{k}W:=W(t_{k+1})-W(t_{k})$, $k\in\mathbb{N}$. We consider the following general one-step approximation of Eq. (1.1):
(2.9) $Y_{k+1}=Y_{k}+B_{b,\sigma}(\tau,Y_{k},Y_{k+1},\delta_{k}W),$
where the explicit or implicit function $B_{b,\sigma}$ depends on the coefficients $b$ and $\sigma$, the step-size $\tau$, the previous value $Y_{k}$, the current value $Y_{k+1}$, and the Brownian increment $\delta_{k}W$ in the subinterval $t_{k},t_{k+1}]$, $k\in\mathbb{N}$.
Denote by $Y_{k}^{x}$ the $k$-th step iterative starting from $Y_{0}=x$. We assume the discrete process $(Y_{k})_{k\in\mathbb{N}}$ is a homogeneous Markov chain with a unique invariant measure $\pi_{\tau}$.
Assume $(Y_{k})_{k\in\mathbb{N}}$ defined in (2.9) has a continuous interpolation $(\hat{Y}_{t})_{t\geq 0}$ which is an Itô process and has the following form in the subinterval $0,\tau]$:
(2.10) $\hat{Y}_{t}=\hat{Y}_{0}+\int_{0}^{t}\hat{b}_{\tau}(\tilde{Y}_{0})\mathrm{d}s+\int_{0}^{t}\hat{\sigma}_{\tau}(\tilde{Y}_{0})\mathrm{d}B_{s},\quad t\in[0,\tau],$
with $\mathscr{F}_{0}$-measurable r.v.s $\hat{Y}_{0}$ and $\tilde{Y}_{0}$ depending on the data of Eq. (1.1). More precisely, we assume that there exists two functions $g_{\tau},\tilde{g}_{\tau}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $\hat{Y}_{0}=g_{\tau}(Y_{0})$ and $\tilde{Y}_{0}=\tilde{g}_{\tau}(Y_{0})$.
Denote by $\hat{Y}_{\tau}^{x}$ the solution $\hat{Y}_{\tau}$ of (2.10) with initial datum $\hat{Y}_{0}=x$ and define a linear operator
(2.11) $\mathcal{A}_{\tau}f(x):=\mathbb{E}[f(\hat{Y}_{\tau}^{x})-f(x)],\quad x\in\mathbb{R}^{d},~f\in\mathcal{C}_{b}^{2}(\mathbb{R}^{d}).$
In particular, when $\hat{Y}_{0}=Y_{0}$ and $\hat{Y}_{\tau}=Y_{1}$, then $\mathcal{A}_{\tau}$ is the generator of the Markov chain $(Y_{k})_{k\in\mathbb{N}}$. Similarly to (2.5), we have
$\mathcal{A}_{\tau}f(x)=$ $\mathbb{E}\int_{0}^{\tau}\Big[\nabla f(\hat{Y}_{s}^{x})\hat{b}_{\tau}(\tilde{g}_{\tau}(x))+\frac{1}{2}\sum_{j=1}^{m}\nabla^{2}f(\hat{Y}_{s}^{x})(\hat{\sigma}_{j,\tau}(\tilde{Y}_{0}),\hat{\sigma}_{j,\tau}(\tilde{g}_{\tau}(x)))\Big]\mathrm{d}s.$
Comparing with (2.4) of $\mathcal{A}f(x)$, we have the following representation of $\mathcal{A}_{\tau}f(x)$:
(2.12) $\mathcal{A}_{\tau}f(y)=\tau\mathcal{A}f(x)+\sum_{i=1}^{6}\mathbb{E}\mathcal{R}_{i}(x,y),$
where
(2.13) $\begin{split}\mathcal{R}_{1}(x,y)&:=\int_{0}^{\tau}(\nabla f(\hat{Y}_{s}^{y})-\nabla f(y))\hat{b}_{\tau}(\tilde{g}_{\tau}(x))\mathrm{d}s,\\ \mathcal{R}_{2}(x,y)&:=\frac{1}{2}\sum_{j=1}^{m}\int_{0}^{\tau}[\nabla^{2}f(\hat{Y}_{s}^{y})-\nabla^{2}f(y)](\hat{\sigma}_{j,\tau}(\tilde{g}_{\tau}(x)),\hat{\sigma}_{j,\tau}(\tilde{g}_{\tau}(x)))\mathrm{d}s,\\ \mathcal{R}_{3}(x,y)&:=(\nabla f(y)-\nabla f(x))\hat{b}_{\tau}(\tilde{g}_{\tau}(x))\tau,\\ \mathcal{R}_{4}(x,y)&:=\frac{1}{2}\sum_{j=1}^{m}[\nabla^{2}f(y)-\nabla^{2}f(x)](\hat{\sigma}_{j,\tau}(\tilde{g}_{\tau}(x)),\hat{\sigma}_{j,\tau}(\tilde{g}_{\tau}(x)))\tau,\\ \mathcal{R}_{5}(x,y)&:=\nabla f(x)[\hat{b}_{\tau}(\tilde{g}_{\tau}(x))-b(x)]\tau,\\ \mathcal{R}_{6}(x,y)&:=\frac{1}{2}\sum_{j=1}^{m}[\nabla^{2}f(x)(\hat{\sigma}_{j,\tau}(\tilde{g}_{\tau}(x)),\hat{\sigma}_{j,\tau}(\tilde{g}_{\tau}(x)))-\nabla^{2}f(x)(\sigma_{j,\tau}(x),\sigma_{j,\tau}(x))]\tau.\end{split}$
with $x,y\in\mathbb{R}^{d}$.
We now state the main result on the ergodic error between the exact invariant measure $\pi$ of Eq. (1.1) and the numerical invariant measure $\pi_{\tau}$ of the associated one-step approximation (2.9).
Theorem 2.3.
For any $\varphi\in\mathcal{B}_{b}(\mathbb{R}^{d})$ or ${\rm Lip}(\mathbb{R}^{d})$ such that $P_{\cdot}\varphi\in L^{1}(0,\infty;\mathcal{C}_{b}^{2}(\mathbb{R}^{d}))$ and $X_{0}\sim\pi_{\tau}$, there holds that
(2.14) $|\pi_{\tau}(\varphi)-\pi(\varphi)|=\tau^{-1}\Big|\mathbb{E}^{\pi_{\tau}}\mathcal{A}_{\tau}f_{\varphi}(g_{\tau}(X_{0}))-\sum_{i=1}^{6}\mathbb{E}^{\pi_{\tau}}[\mathbb{E}\mathcal{R}_{i}(x,g_{\tau}(x))|_{x=X_{0}}]\Big|.$
Proof.
Let $X_{0}\sim\pi_{\tau}$ and $\varphi\in\mathcal{B}_{b}(\mathbb{R}^{d})$ or ${\rm Lip}(\mathbb{R}^{d})$ such that $P_{\cdot}\varphi\in L^{1}(0,\infty;\mathcal{C}_{b}^{2}(\mathbb{R}^{d}))$. It follows from the Stein equation (2.6) and the representation (2.12), we have
$|\pi_{\tau}(\varphi)-\pi(\varphi)|=$ $|\mathbb{E}^{\pi_{\tau}}\varphi(X_{0})-\pi(\varphi)|$
$=$ $|\mathbb{E}^{\pi_{\tau}}\mathcal{A}f_{\varphi}(X_{0})|$
$=$ $\tau^{-1}\Big|\mathbb{E}^{\pi_{\tau}}\mathcal{A}_{\tau}f_{\varphi}(g_{\tau}(X_{0}))-\sum_{i=1}^{6}\mathbb{E}^{\pi_{\tau}}[\mathbb{E}\mathcal{R}_{i}(x,g_{\tau}(x))|_{x=X_{0}}]\Big|,$
which proves (2.14). ∎
3. Applications to Superlinear SODEs
This section will apply the results in Theorem 2.3 to a family of one-step approximations for superlinear SODEs.
Though one can utilize the same methodology to handle a large class of examples, we focus on several recently studied one-step numerical schemes–including the tamed Euler, projected Euler, and backward Euler methods in Example 3.3.
3.1. Assumptions and Preliminary Estimates
We impose the following frequently used dissipativity and polynomial growth conditions.
Assumption 3.1.
There exist positive constants $L_{1},L_{2},L_{3}$, $p^{\star}\geq 2$, and $\gamma>1$ such that for any $x,y\in\mathbb{R}^{d}$,
(3.1) $\langle x-y,b(x)-b(y)\rangle+\frac{2p^{\star}-1}{2}\|\sigma(x)-\sigma(y)\|^{2}\leq-L_{1}|x-y|^{2},$
(3.2) $\langle x,b(x)\rangle+\frac{p^{\star}(2p^{\star}-1)}{2}\|\sigma(x)\|_{HS}^{2}\leq L_{2}-L_{3}|x|^{\gamma+1}.$
Assumption 3.2.
The functions $b,\sigma$ have continuous derivatives up to order four, and there exists $C>0$ such that for any $k=1,2,3,4$ and $x,v_{k}\in\mathbb{R}^{d}$,
$|\nabla^{k}b(x)(v_{1},\cdots,v_{k})|$ $\leq C\big[\mathbf{1}_{\gamma\leq k}+\mathbf{1}_{\gamma>k}(1+|x|^{\gamma-k})\big]|v_{1}|\cdots|v_{k}|,$
$|\nabla^{k}\sigma_{j}(x)(v_{1},\cdots,v_{k})|^{2}$ $\leq C[\mathbf{1}_{\gamma\leq 2k-1}+\mathbf{1}_{\gamma>2k-1}(1+|x|^{\gamma-(2k-1)})]|v_{1}|^{2}\cdots|v_{k}|^{2}.$
Remark 3.1.
Under Assumption 3.1, Eq. (1.1) has a unique strong solution, which is a Markov process exponentially ergodic to a unique invariant measure $\pi$ [8, Theorem 3.1.1, Proposition 4.3.5, and Theorem 4.3.9]. Moreover, if $X_{0}\in L^{2p^{*}}$, then
$\sup_{t\geq 0}\mathbb{E}|X_{t}|^{2p^{*}}\leq C(1+\mathbb{E}|X_{0}|^{2p^{*}}),$
provided $X_{0}\in L^{2p^{*}}$. Under the additional Assumption 3.2, it is clear, for all $T>0$ and $j=\{1,2,\cdots,m\}$, that $\mathbb{E}\int_{0}^{T}|\sigma_{j}(X_{t})|^{2}\mathrm{d}t\leq CT(1+\mathbb{E}|X_{0}|^{\gamma+1})<\infty$ so that $\int_{0}^{\cdot}\sigma(X_{s})\mathrm{d}W(s)$ is a martingale; see, e.g., [9].
3.2. Estimates of Stein Equation
To apply Theorems 2.1 and 2.3 in the present settings, we need to show that $P_{\cdot}\varphi\in L^{1}(0,\infty;\mathcal{C}_{b}^{2}(\mathbb{R}^{d}))$ for a given test function $\varphi$ and to derive necessary derivative estimates for the Stein equation solution $f_{\varphi}$.
We have the following estimates on the directional derivative $\nabla^{i}X_{t}^{\cdot}$ with $i=1,2,3,4$ and $t\geq 0$; see [12, Lemma 3.3] for a weak result in the case of weak dissipativity.
Proposition 3.2.
Let Assumptions 3.1 and 3.2 hold and $q\in(1,p^{\star})$. There exist positive constants $\lambda$, $\gamma_{1}^{*}$, and $C$ such that
(3.3) $\sum_{i=1}^{4}\|\nabla^{i}X_{t}^{x}\|_{L^{2q}(\Omega;\mathcal{L}((\mathbb{R}^{d})^{\otimes i};\mathbb{R}))}$ $\leq Ce^{-\lambda t}(1+|x|^{\gamma_{1}^{*}}),\quad x\in\mathbb{R}^{d},~t\geq 0.$
Consequently, $P_{\cdot}\varphi\in L^{1}(0,\infty;\mathcal{C}_{b}^{2}(\mathbb{R}^{d}))$ for any $\varphi\in\mathcal{C}_{b}^{2}(\mathbb{R}^{d})$ and there exist positive constant $C$ and $\gamma_{1}$ such that for any $\varphi\in\mathcal{C}_{b}^{i}(\mathbb{R}^{d})$,
(3.4) $\|\nabla^{i}f_{\varphi}^{x}\|_{\mathcal{L}((\mathbb{R}^{d})^{\otimes i};\mathbb{R})}$ $\leq C(1+|x|^{\gamma_{1}})\|\nabla^{i}\varphi\|,\quad i=1,2,3,4.$
Proof.
Let $q\in(1,p^{\star})$, $t\geq 0$, and $x,v_{1},v_{2},v_{3},v_{4}\in\mathbb{R}^{d}$. In what follows, we denote
$\eta^{v_{1}}(t,x):=\nabla X_{t}^{x}v_{1},\quad$ $\xi^{v_{1},v_{2}}(t,x):=\nabla^{2}X_{t}^{x}(v_{1},v_{2}),$
$\zeta^{v_{1},v_{2},v_{3}}(t,x):=\nabla^{3}X_{t}^{x}(v_{1},v_{2},v_{3}),\quad$ $\varpi^{v_{1},v_{2},v_{3},v_{4}}(t,x):=\nabla^{4}X_{t}^{x}(v_{1},v_{2},v_{3},v_{4}).$
It follows from (3.1) that
(3.5) $\langle\nabla b(x)y,y\rangle+\frac{2p^{\star}-1}{2}\|D\sigma(x)y\|^{2}\leq-L_{1}|y|^{2},\quad x,y\in\mathbb{R}^{d}.$
For $\eta^{v_{1}}(t,x)$, we have $\eta^{v_{1}}(0,x)=v_{1}$ and for $t>0$,
$\mathrm{d}\eta^{v_{1}}(t,x)=\nabla b(X_{t}^{x})\eta^{v_{1}}(t,x)\mathrm{d}t+\sum_{j=1}^{m}\nabla\sigma_{j}(X_{t}^{x})\eta^{v_{1}}(t,x)\mathrm{d}W_{j}(t).$
Define a stopping time
$\tau_{n}:=\inf\{s\geq 0:|\eta^{v_{1}}(s,x)|>n\}.$
Using Itô formula and the inequality (3.5) implies
$\frac{1}{2q}[\mathbb{E}|\eta^{v_{1}}(t\wedge\tau_{n},x)|^{2q}-|v_{1}|^{2q}]$
$=\mathbb{E}\int_{0}^{t\wedge\tau_{n}}|\eta^{v_{1}}(x)|^{2(q-1)}\Big[\langle\eta^{v_{1}}(x),\nabla b(X^{x})\eta^{v_{1}}(x)\rangle$
$\qquad+\frac{2q-1}{2}\sum_{j=1}^{m}|\nabla\sigma_{j}(X^{x})\eta^{v_{1}}(x)|^{2}\Big]\mathrm{d}s$
$\leq-L_{1}\mathbb{E}\int_{0}^{t\wedge\tau_{n}}|\eta^{v_{1}}(x)|^{2q}\mathrm{d}s.$
We omit the integration variable here and after when there is an integration to lighten the notation. Using Gronwall inequality and Fatou lemma, we get
(3.6) $\mathbb{E}|\nabla X_{t}^{x}v_{1}|^{2q}$ $\leq e^{-2q\lambda t}|v_{1}|^{2q}.$
Similarly, $\xi^{v_{1},v_{2}}(t,x)$ satisfies $\xi^{v_{1},v_{2}}(0,x)=0$ and
$\mathrm{d}\xi^{v_{1},v_{2}}(t,x)$ $=[\nabla b(X_{t}^{x})\xi^{v_{1},v_{2}}(t,x)+\nabla^{2}b(X_{t}^{x})(\eta^{v_{1}}(t,x),\eta^{v_{2}}(t,x))]\mathrm{d}t$
$\quad+\sum_{j=1}^{m}[\nabla\sigma_{j}(X_{t}^{x})\xi^{v_{1},v_{2}}(t,x)+\nabla^{2}\sigma_{j}(X_{t}^{x})(\eta^{v_{1}}(t,x),\eta^{v_{2}}(t,x))]\mathrm{d}W_{j}(t),$
so that (introducing an analogous stopping time if necessary)
$\frac{1}{2q}\mathbb{E}|\xi^{v_{1},v_{2}}(t,x)|^{2q}$
$=\mathbb{E}\int_{0}^{t}|\xi^{v_{1},v_{2}}(x)|^{2(q-1)}\Big[\langle\xi^{v_{1},v_{2}}(x),\nabla b(X^{x})\xi^{v_{1},v_{2}}(x)+\nabla^{2}b(X^{x})(\eta^{v_{1}}(x),\eta^{v_{2}}(x))\rangle$
$\qquad+\frac{2q-1}{2}\sum_{j=1}^{m}|\nabla\sigma_{j}(X^{x})\xi^{v_{1},v_{2}}(x)+\nabla^{2}\sigma_{j}(X^{x})(\eta^{v_{1}}(x),\eta^{v_{2}}(x))|^{2}\Big]\mathrm{d}s.$
By Young inequality and (3.5), for an arbitrarily small $\epsilon>0$ such that $\frac{2q-1}{2}(1+\epsilon)\leq\frac{2p^{\star}-1}{2}$, there exists $\delta=\delta(\epsilon)\in(0,L_{1})$ and $C_{\epsilon,\delta}>0$ such that
$\frac{1}{2q}\mathbb{E}|\xi^{v_{1},v_{2}}(t,x)|^{2q}$
$\leq\mathbb{E}\int_{0}^{t}|\xi^{v_{1},v_{2}}(x)|^{2(q-1)}\big[\epsilon|\xi^{v_{1},v_{2}}(x)|^{2}+(2\epsilon)^{-1}|\nabla^{2}b(X^{x})(\eta^{v_{1}}(x),\eta^{v_{2}}(x))|^{2}\big]\mathrm{d}s$
$\quad+\mathbb{E}\int_{0}^{t}|\xi^{v_{1},v_{2}}(x)|^{2(q-1)}[\langle\xi^{v_{1},v_{2}}(x),\nabla b(X^{x})\xi^{v_{1},v_{2}}(x)\rangle$
$\qquad\qquad\qquad\qquad\qquad\qquad+\frac{2q-1}{2}(1+\epsilon)\sum_{j=1}^{m}|\nabla\sigma_{j}(X^{x})\xi^{v_{1},v_{2}}(x)|^{2}]\mathrm{d}s$
$\quad+\frac{2q-1}{2}(1+\epsilon^{-1})\sum_{j=1}^{m}\mathbb{E}\int_{0}^{t}|\xi^{v_{1},v_{2}}(x)|^{2(q-1)}|\nabla^{2}\sigma_{j}(X^{x})(\eta^{v_{1}}(x),\eta^{v_{2}}(x))|^{2}\mathrm{d}s$
$\leq-(L_{1}-\delta)\int_{0}^{t}\mathbb{E}|\xi^{v_{1},v_{2}}(x)|^{2q}\mathrm{d}s+C_{\epsilon,\delta}\underbrace{\mathbb{E}\int_{0}^{t}|\nabla^{2}b(X^{x})(\eta^{v_{1}}(x),\eta^{v_{2}}(x))|^{2q}\mathrm{d}s}_{I_{11}}$
$\quad+C_{\epsilon,\delta}\underbrace{\sum_{j=1}^{m}\mathbb{E}\int_{0}^{t}|\nabla^{2}\sigma_{j}(X^{x})(\eta^{v_{1}}(x),\eta^{v_{2}}(x))|^{2q}\mathrm{d}s}_{I_{12}}.$
It remains to estimate $I_{11}$ and $I_{12}$. We employ Assumption 3.2, Hölder inequality, and the estimate (3.6) to arrive at
$I_{11}\leq$ $C\mathbb{E}\int_{0}^{t}\big[\mathbf{1}_{\gamma\in(1,2]}+\mathbf{1}_{\gamma>2}(1+|X^{x}|^{\gamma-2})\big]^{2q}|\eta^{v_{1}}(x)|^{2q}|\eta^{v_{2}}(x)|^{2q}\mathrm{d}s$
$\leq$ $C\big[\mathbf{1}_{\gamma\in(1,2]}+\mathbf{1}_{\gamma>2}(1+|x|^{2q(\gamma-2)})\big]|v_{1}|^{2q}|v_{2}|^{2q}.$
In the same manner,
$I_{12}\leq C\big[\mathbf{1}_{\gamma\in(1,3]}+\mathbf{1}_{\gamma>3}(1+|x|^{q(\gamma-3)})\big]|v_{1}|^{2q}|v_{2}|^{2q}.$
Combining the above two estimates, we have
$\frac{1}{2q}\mathbb{E}|\xi^{v_{1},v_{2}}(t,x)|^{2q}+(L_{1}-\delta)\mathbb{E}\int_{0}^{t}|\xi^{v_{1},v_{2}}(x)|^{2q}\mathrm{d}s$
$\leq C[\mathbf{1}_{\gamma\in(1,2]}+\mathbf{1}_{\gamma\in[2,3]}(1+|x|^{2q(\gamma-2)})+\mathbf{1}_{\gamma>3}(1+|x|^{q(\gamma-3)})]|v_{1}|^{2q}|v_{2}|^{2q}.$
and in combination with Gronwall inequality, we derive
(3.7) $\mathbb{E}|\nabla^{2}X_{t}^{x}(v_{1},v_{2})|^{2q}$ $\leq Ce^{-2q\lambda t}[\mathbf{1}_{\gamma\in(1,2]}+\mathbf{1}_{\gamma>2}(1+|x|^{2q(\gamma-2)})]|v_{1}|^{2q}|v_{2}|^{2q}.$
For the third term $\zeta^{v_{1},v_{2},v_{3}}(t,{x})$, we get $\zeta^{v_{1},v_{2},v_{3}}(0,x)=0$ and
$\zeta^{v_{1},v_{2},v_{3}}(t,x)$
$=(\nabla b(X_{t}^{x})\zeta^{v_{1},v_{2},v_{3}}(t,x)+\nabla^{2}b(X_{t}^{x})(\eta^{v_{1}}(t,x),\xi^{v_{2},v_{3}}(t,x))$
$\quad+\nabla^{2}b(X_{t}^{x})(\xi^{v_{1},v_{3}}(t,x),\eta^{v_{2}}(t,x))+\nabla^{2}b(X_{t}^{x})(\xi^{v_{1},v_{2}}(t,x),\eta^{v_{3}}(t,x))$
$\quad+\nabla^{3}b(X_{t}^{x})(\eta^{v_{1}}(t,x),\eta^{v_{2}}(t,x),\eta^{v_{3}}(t,x)))\mathrm{d}t$
$\quad+\sum_{j=1}^{m}(\nabla\sigma_{j}(X_{t}^{x})\zeta^{v_{1},v_{2},v_{3}}(t,x)+\nabla^{2}\sigma_{j}(X_{t}^{x})(\eta^{v_{1}}(t,x),\xi^{v_{2},v_{3}}(t,x))$
$\quad+\nabla^{2}\sigma_{j}(X_{t}^{x})(\xi^{v_{1},v_{3}}(t,x),\eta^{v_{2}}(t,x))+\nabla^{2}\sigma_{j}(X_{t}^{x})(\xi^{v_{2},v_{3}}(t,x),\eta^{v_{1}}(t,x))$
$\quad+\nabla^{3}\sigma_{j}(X_{t}^{x})(\eta^{v_{1}}(t,x),\eta^{v_{2}}(t,x),\eta^{v_{3}}(t,x)))\mathrm{d}W_{j}(t)$
$=:(\nabla b(X_{t}^{x})\zeta^{v_{1},v_{2},v_{3}}(t,x)+H(X_{t}^{x}))\mathrm{d}t$
$\quad+\sum_{j=1}^{m}(\nabla\sigma_{j}(X_{t}^{x})\zeta^{v_{1},v_{2},v_{3}}(t,x)+G_{j}(X_{t}^{x}))\mathrm{d}W_{j}(t).$
As in the same arguments to deal with the term $\xi^{v_{1},v_{2}}(t,x)$, we have
$\quad\frac{1}{2q}\mathbb{E}|\zeta^{v_{1},v_{2},v_{3}}(t,x)|^{2q}+(L_{1}-\delta)\mathbb{E}\int_{0}^{t}|\zeta^{v_{1},v_{2},v_{3}}(x)|^{2q}\mathrm{d}s$
$\leq C\underbrace{\mathbb{E}\int_{0}^{t}|H(X^{x})|^{2q}\mathrm{d}s}_{I_{21}}+C\underbrace{\sum_{j=1}^{m}\mathbb{E}\int_{0}^{t}|G_{j}(X^{x})|^{2q}\mathrm{d}s}_{I_{22}}.$
We split the estimation of the term $I_{21}$ into the following four parts:
$I_{21}$ $\leq C\int_{0}^{t}\mathbb{E}|\nabla^{2}b(X^{x})(\eta^{v_{1}}(x),\xi^{v_{2},v_{3}}(x))|^{2q}\mathrm{d}s$
$\quad+C\int_{0}^{t}\mathbb{E}|\nabla^{2}b(X^{x})(\xi^{v_{1},v_{3}}(x),\eta^{v_{2}}(x))|^{2q}\mathrm{d}s$
$\quad+C\int_{0}^{t}\mathbb{E}|\nabla^{2}b(X^{x})(\xi^{v_{1},v_{2}}(x),\eta^{v_{3}}(x))|^{2q}\mathrm{d}s$
$\quad+C\int_{0}^{t}\mathbb{E}|\nabla^{3}b(X^{x})(\eta^{v_{1}}(x),\eta^{v_{2}}(x),\eta^{v_{3}}(x))|^{2q}\mathrm{d}s$
$=:I_{211}+I_{212}+I_{213}+I_{214}.$
We mention that the analysis of $I_{211},I_{212},I_{213}$ is similar and we take $I_{211}$ as an example. Bt Assumption 3.2, Hölder inequality, and the estimates (3.6) and (3.7), we have
$I_{211}\leq C(1+|x|^{4q(\gamma-2)})|v_{1}|^{2q}|v_{2}|^{2q}.$
For $I_{214}$, we have
$I_{214}\leq C[\mathbf{1}_{\gamma\in(1,3]}+\mathbf{1}_{\gamma>3}(1+|x|^{2q(\gamma-3)})]|v_{1}|^{2q}|v_{2}|^{2q}|v_{3}|^{2q}.$
Hence, we get
$I_{21}$ $\leq C[\mathbf{1}_{1<\gamma\leq 2}+\mathbf{1}_{\gamma>2}(1+|x|^{4q(\gamma-2)})]|v_{1}|^{2q}|v_{2}|^{2q}|v_{3}|^{2q}.$
Similarly, it is not hard to obtain
$I_{22}$ $\leq C[\mathbf{1}_{\gamma\in(1,2]}+\mathbf{1}_{\gamma>2}(1+|x|^{4q(\gamma-2)})]|v_{1}|^{2q}|v_{2}|^{2q}|v_{3}|^{2q}.$
Plugging these estimates together yields
$\frac{1}{2q}\mathbb{E}[|\zeta^{v_{1},v_{2},v_{3}}(t,x)|^{2q}+(L_{1}-\delta)\mathbb{E}\int_{0}^{t}|\zeta^{v_{1},v_{2},v_{3}}(x)|^{2q}\mathrm{d}s$
$\leq C[\mathbf{1}_{1<\gamma\leq 2}+\mathbf{1}_{\gamma>2}(1+|x|^{4q(\gamma-2)})]|v_{1}|^{2q}|v_{2}|^{2q}|v_{3}|^{2q}.$
By Gronwall inequality, we conclude
(3.8) $\mathbb{E}|\nabla^{3}X_{t}^{x}(v_{1},v_{2},v_{3})|^{2q}$ $\leq Ce^{-2q\lambda t}[\mathbf{1}_{1<\gamma\leq 2}+\mathbf{1}_{\gamma>2}(1+|x|^{4q(\gamma-2)})]|v_{1}|^{2q}|v_{2}|^{2q}|v_{3}|^{2q}.$
For the last term, we get $\varpi^{v_{1},v_{2},v_{3},v_{4}}(0,x)=0$ and
$\varpi^{v_{1},v_{2},v_{3},v_{4}}(t,x)=$ $(\nabla b(X_{t}^{x})\varpi^{v_{1},v_{2},v_{3},v_{4}}(t,x)+J_{31})\mathrm{d}t$
$+\sum_{j=1}^{m}(\nabla\sigma_{j}(X_{t}^{x})\varpi^{v_{1},v_{2},v_{3},v_{4}}(t,x)+J^{j}_{32})\mathrm{d}W_{j}(t).$
where
$J_{31}=$ $\sum_{i=1}^{14}J_{31i}$
$=$ $\nabla^{2}b(X_{t}^{x})(\eta^{v_{1}}(t,x),\zeta^{v_{2},v_{3},v_{4}}(t,x))$
$+\nabla^{2}b(X_{t}^{x})(\eta^{v_{2}}(t,x),\zeta^{v_{1},v_{3},v_{4}}(t,x))$
$+\nabla^{2}b(X_{t}^{x})(\eta^{v_{3}}(t,x),\zeta^{v_{1},v_{2},v_{4}}(t,x))$
$+\nabla^{2}b(X_{t}^{x})(\eta^{v_{4}}(t,x),\zeta^{v_{1},v_{2},v_{3}}(t,x))$
$+\nabla^{2}b(X_{t}^{x})(\xi^{v_{1},v_{2}}(t,x),\xi^{v_{3},v_{4}}(t,x))$
$+\nabla^{2}b(X_{t}^{x})(\xi^{v_{1},v_{3}}(t,x),\xi^{v_{2},v_{4}}(t,x))$
$+\nabla^{2}b(X_{t}^{x})(\xi^{v_{1},v_{4}}(t,x),\xi^{v_{2},v_{3}}(t,x))$
$+\nabla^{3}b(X_{t}^{x})(\xi^{v_{1},v_{2}}(t,x),\eta^{v_{3}}(t,x),\eta^{v_{4}}(t,x))$
$+\nabla^{3}b(X_{t}^{x})(\xi^{v_{1},v_{3}}(t,x),\eta^{v_{2}}(t,x),\eta^{v_{4}}(t,x))$
$+\nabla^{3}b(X_{t}^{x})(\xi^{v_{1},v_{4}}(t,x),\eta^{v_{2}}(t,x),\eta^{v_{3}}(t,x))$
$+\nabla^{3}b(X_{t}^{x})(\xi^{v_{2},v_{3}}(t,x),\eta^{v_{1}}(t,x),\eta^{v_{4}}(t,x))$
$+\nabla^{3}b(X_{t}^{x})(\xi^{v_{2},v_{4}}(t,x),\eta^{v_{1}}(t,x),\eta^{v_{3}}(t,x))$
$+\nabla^{3}b(X_{t}^{x})(\xi^{v_{3},v_{4}}(t,x),\eta^{v_{1}}(t,x),\eta^{v_{2}}(t,x))$
$+\nabla^{4}b(X_{t}^{x})(\eta^{v_{1}}(t,x),\eta^{v_{2}}(t,x),\eta^{v_{3}}(t,x),\eta^{v_{4}}(t,x)),$
$J^{j}_{32}=$ $\sum_{i=1}^{14}J^{j}_{32i}$
$=$ $\nabla^{2}\sigma_{j}(X_{t}^{x})(\eta^{v_{1}}(t,x),\zeta^{v_{2},v_{3},v_{4}}(t,x))$
$+\nabla^{2}\sigma_{j}(X_{t}^{x})(\eta^{v_{2}}(t,x),\zeta^{v_{1},v_{3},v_{4}}(t,x))$
$+\nabla^{2}\sigma_{j}(X_{t}^{x})(\eta^{v_{3}}(t,x),\zeta^{v_{1},v_{2},v_{4}}(t,x))$
$+\nabla^{2}\sigma_{j}(X_{t}^{x})(\eta^{v_{4}}(t,x),\zeta^{v_{1},v_{2},v_{3}}(t,x))$
$+\nabla^{2}\sigma_{j}(X_{t}^{x})(\xi^{v_{1},v_{2}}(t,x),\xi^{v_{3},v_{4}}(t,x))$
$+\nabla^{2}\sigma_{j}(X_{t}^{x})(\xi^{v_{1},v_{3}}(t,x),\xi^{v_{2},v_{4}}(t,x))$
$+\nabla^{2}\sigma_{j}(X_{t}^{x})(\xi^{v_{1},v_{4}}(t,x),\xi^{v_{2},v_{3}}(t,x))$
$+\nabla^{3}\sigma_{j}(X_{t}^{x})(\xi^{v_{1},v_{2}}(t,x),\eta^{v_{3}}(x,t),\eta^{v_{4}}(t,x))$
$+\nabla^{3}\sigma_{j}(X_{t}^{x})(\xi^{v_{1},v_{3}}(t,x),\eta^{v_{2}}(t,x),\eta^{v_{4}}(t,x))$
$+\nabla^{3}\sigma_{j}(X_{t}^{x})(\xi^{v_{1},v_{4}}(t,x),\eta^{v_{2}}(t,x),\eta^{v_{3}}(t,x))$
$+\nabla^{3}\sigma_{j}(X_{t}^{x})(\xi^{v_{2},v_{3}}(t,x),\eta^{v_{1}}(t,x),\eta^{v_{4}}(t,x))$
$+\nabla^{3}\sigma_{j}(X_{t}^{x})(\xi^{v_{2},v_{4}}(t,x),\eta^{v_{1}}(t,x),\eta^{v_{3}}(t,x))$
$+\nabla^{3}\sigma_{j}(X_{t}^{x})(\xi^{v_{3},v_{4}}(t,x),\eta^{v_{1}}(t,x),\eta^{v_{2}}(t,x))$
$+\nabla^{4}\sigma_{j}(X_{t}^{x})(\eta^{v_{1}}(t,x),\eta^{v_{2}}(t,x),\eta^{v_{3}}(t,x),\eta^{v_{4}}(t,x)).$
Then we have
$\frac{1}{2q}\mathbb{E}|\varpi^{v_{1},v_{2},v_{3},v_{4}}(t,x)|^{2q}+(L_{1}-\delta)\mathbb{E}\int_{0}^{t}|\varpi^{v_{1},v_{2},v_{3},v_{4}}(x)|^{2q}\mathrm{d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}|J_{31}|^{2q}\mathrm{d}s+C\sum_{j=1}^{m}\mathbb{E}\int_{0}^{t}|J^{j}_{32}|^{2q}\mathrm{d}s$
$\leq$ $C\sum_{i=1}^{14}\mathbb{E}\int_{0}^{t}|J_{31i}|^{2q}\mathrm{d}s+C\sum_{i=1}^{14}\sum_{j=1}^{m}\mathbb{E}\int_{0}^{t}|J^{j}_{32i}|^{2q}\mathrm{d}s.$
For $\sum_{i=1}^{4}\mathbb{E}\int_{0}^{t}|J_{31i}|^{2q}\mathrm{d}s$, we employ Assumption (3.2), Hölder inequality, the estimates (3.6), (3.7), and (3.8) to get
$\sum_{i=1}^{14}\mathbb{E}\int_{0}^{t}|J_{31i}|^{2q}\mathrm{d}s\leq C[\mathbf{1}_{\gamma\in(1,2]}+\mathbf{1}_{\gamma\geq 2}(1+|x|^{6q(\gamma-2)})]|v_{1}|^{2q}|v_{2}|^{2q}|v_{3}|^{2q}|v_{4}|^{2q}.$
Similarly, we have
$\sum_{i=1}^{14}\sum_{j=1}^{m}\mathbb{E}\int_{0}^{t}|J^{j}_{31i}|^{2q}\mathrm{d}s\leq C[\mathbf{1}_{\gamma\in(1,2]}+\mathbf{1}_{\gamma\geq 2}(1+|x|^{6q(\gamma-2)})]|v_{1}|^{2q}|v_{2}|^{2q}|v_{3}|^{2q}|v_{4}|^{2q}.$
Hence
$\frac{1}{2q}\mathbb{E}|\varpi^{v_{1},v_{2},v_{3},v_{4}}(t,x)|^{2q}+(L_{1}-\delta)\mathbb{E}\int_{0}^{t}|\varpi^{v_{1},v_{2},v_{3},v_{4}}(x)|^{2q}\mathrm{d}s$
$\leq C[\mathbf{1}_{\gamma\in(1,2]}+\mathbf{1}_{\gamma\geq 2}(1+|x|^{6q(\gamma-2)})]|v_{1}|^{2q}|v_{2}|^{2q}|v_{3}|^{2q}|v_{4}|^{2q},$
and we conclude
$\mathbb{E}|\nabla^{4}X_{t}^{x}(v_{1},v_{2},v_{3},v_{4})|^{2q}$
(3.9) $\leq Ce^{-2q\lambda t}[\mathbf{1}_{\gamma\in(1,2]}+\mathbf{1}_{\gamma>2}(1+|x|^{6q(\gamma-2)})]|v_{1}|^{2q}|v_{2}|^{2q}|v_{3}|^{2q}|v_{4}|^{2q}.$
Thus, we conclude (3.3) by the estimates (3.6), (3.7), (3.8), and (3.2).
Finally, it follows from (2.7), (2.1), and (3.3) with $i=1$ that
$\|\nabla f_{\varphi}(x)\|_{\mathcal{L}(\mathbb{R}^{d})}$ $\leq\|\nabla\varphi\|_{1}\int_{0}^{+\infty}\mathbb{E}\|\nabla X_{t}^{x}\|_{\mathcal{L}(\mathbb{R}^{d})}\mathrm{d}t\leq C_{1}(1+|x|^{\gamma_{1}})\|\nabla\varphi\|_{1},$
which shows the estimate (3.4) with $i=1$. The remaining estimates (3.4) with $i=2,3,4$ are analogously, and we omit the details. ∎
3.3. Ergodic Estimates of Numerical Schemes
We apply Theorem 2.3 to this part’s implicit and explicit numerical schemes.
Example 3.3.
(1) MEM:
(3.10) $Y_{n+1}=\mathscr{P}(Y_{n})+b_{\tau}(\mathscr{P}(Y_{n}))\tau+\sigma_{\tau}(\mathscr{P}(Y_{n}))\delta_{n}W,\quad n\in\mathbb{N},$
where $\mathscr{P}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is a modification function, and $b_{\tau}(x)$ and $\sigma_{j,\tau}(x)$, $j\in\{1,2,\cdots,m\}$, are $\mathbb{R}^{d}$-valued measurable functions. This includes the following choices.
(a) TEM:
(3.11) $\begin{split}\mathscr{P}(x):=x,~b_{\tau}(x):=\frac{b(x)}{(1+\tau|x|^{4(\gamma-1)})^{1/4}},\quad\sigma_{\tau}(x):=\frac{\sigma(x)}{(1+\tau|x|^{4(\gamma-1)})^{1/4}},\quad x\in\mathbb{R}^{d}.\end{split}$
(b) PEM:
(3.12) $\begin{split}\mathscr{P}(x)&:=\left\{\begin{array}[]{ll}\min\{1,\tau^{-\frac{1}{2\gamma}}|x|^{-1}\}x,&x\neq 0,\\ 0,&x=0,\end{array}\right.~b_{\tau}(x):=b(x),~\sigma_{\tau}(x):=\sigma(x),\quad x\in\mathbb{R}^{d}.\end{split}$
(2) BEM:
(3.13) $Y_{n+1}=Y_{n}+b(Y_{n+1})\tau+\sigma(Y_{n})\delta_{n}W,\quad n\in\mathbb{N}.$
Remark 3.4.
(1) The discrete processes $(Y_{k})_{k\in\mathbb{N}}$ given in Example 3.3 are homogeneous Markov chains and possess unique invariant measures, denoted by $\pi^{M}_{\tau}$ (for MEM) and $\pi^{B}_{\tau}$ (for BEM), respectively [12, 11].
(2) The following estimates hold [12, Example 2.2]: there exist positive constants $C$, $\alpha_{1}\geq 1$, and $\alpha_{2}\geq 1$ such that for any $x\in\mathbb{R}^{d}$ and $j\in\{1,2,\cdots,m\}$,
(3.14) $\begin{split}|\mathscr{P}(x)|&\leq|x|,\\ |b_{\tau}(\mathscr{P}(x))|&\leq C|b(x)|,\\ |\sigma_{j,\tau}(\mathscr{P}(x))|&\leq C|\sigma_{j}(x)|,\\ |b_{\tau}(\mathscr{P}(x))-b(\mathscr{P}(x))|+\sum_{j=1}^{m}|\sigma_{j,\tau}(\mathscr{P}(x))-\sigma_{j}(\mathscr{P}(x))|&\leq C\tau(1+|x|^{\alpha_{1}}),\\ |\mathscr{P}(x)-x|&\leq C\tau^{2}|x|^{\alpha_{2}}.\end{split}$
For TEM (3.11), $\alpha_{1}=5\gamma-4$ and $\alpha_{2}>0$ arbitrary; for PEM (3.12), $\alpha_{1}=1$ and $\alpha_{2}=4\gamma+1$.
Theorem 3.5.
Let Assumptions 3.1 and 3.2 hold with $2p^{\star}\geq\max\{5\gamma-4,4\gamma+1\}$. There exists a positive constant $C$ such that
(3.15) $\sup_{\|\varphi\|_{4}\leq 1}|\pi_{\tau}^{M}(\varphi)-\pi(\varphi)|$ $\leq C\tau,$
(3.16) $\sup_{\|\varphi\|_{4}\leq 1}|\pi_{\tau}^{B}(\varphi)-\pi(\varphi)|$ $\leq C\tau.$
Proof.
Let $\|\varphi\|_{4}\leq 1$.
(1) For the MEM (3.10), the associated intepolation process $(\hat{Y}_{t})_{t\in[0,\tau]}$ can be choosen as
(3.17) $\hat{Y}_{t}=\mathscr{P}(Y_{0})+\int_{0}^{t}b_{\tau}(\mathscr{P}(Y_{0}))\mathrm{d}s+\int_{0}^{t}\sigma_{\tau}(\mathscr{P}(Y_{0}))\mathrm{d}W(s),\quad t\in[0,\tau],$
i.e., $g_{\tau}(x)=\tilde{g}_{\tau}(x)=\mathscr{P}(x)$, $x\in\mathbb{R}^{d}$. It is straightforward to verify that $\hat{Y}_{\tau}=Y_{1}$, and that $\hat{Y}$ is a continuous Itô process.
Let us first estimate the term $\mathbb{E}^{\pi_{\tau}}\mathcal{A}_{\tau}f_{\varphi}(\mathscr{P}(Y_{0}))$ with $Y_{0}=X_{0}\sim\pi_{\tau}^{M}$. As $\pi_{\tau}^{M}$ is an invariant measure of (3.10), we have $\hat{Y}_{\tau}=Y_{1}\sim\pi_{\tau}^{M}$. Thus, $f(\hat{Y}_{\tau})$ and $f(Y_{0})$ share the same law so that $\mathbb{E}^{\pi_{\tau}}[\mathbb{E}[f(\hat{Y}_{\tau}^{x})-f(Y_{0})]|_{x=\mathscr{P}(Y_{0})}]=0$. Then (2.11) implies
$|\mathbb{E}^{\pi_{\tau}}\mathcal{A}_{\tau}f_{\varphi}(\mathscr{P}(Y_{0}))|$ $=|\mathbb{E}^{\pi_{\tau}}\mathbb{E}[f(Y_{0})-f(x)]|_{x=\mathscr{P}(Y_{0})}]|$
(3.18) $\leq\|\nabla f\|\mathbb{E}^{\pi_{\tau}}|Y_{0}-\mathscr{P}(Y_{0})|.$
Let $y=\mathscr{P}(x)$. By the first two identities of the representation (2.13), Itô formula, and the identity (2.5), we have
$|\mathbb{E}\mathcal{R}_{1}(x,g_{\tau}(x))|$
$=\Big|\int_{0}^{\tau}\mathbb{E}\int_{0}^{s}\Big[\nabla^{2}f(\hat{Y}_{r}^{y})(b_{\tau}(\mathscr{P}(x)),b_{\tau}(\mathscr{P}(x)))$
$\quad+\frac{1}{2}\sum_{j=1}^{m}\nabla^{3}f(\hat{Y}_{r}^{y})(b_{\tau}(\mathscr{P}(x)),\sigma_{j,\tau}(\mathscr{P}(x)),\sigma_{j,\tau}(\mathscr{P}(x)))\Big]\mathrm{d}r\mathrm{d}s\Big|$
$\leq\frac{1}{4}(2\|\nabla^{2}f\||b_{\tau}(\mathscr{P}(x))|^{2}+\|\nabla^{3}f\|\|\sigma_{\tau}(\mathscr{P}(x))\|_{HS}^{2}|b_{\tau}(\mathscr{P}(x))|)\tau^{2},$
$|\mathbb{E}\mathcal{R}_{2}(x,g_{\tau}(x))|$
$=\frac{1}{2}\Big|\int_{0}^{\tau}\sum_{j=1}^{m}\mathbb{E}\Big[\int_{0}^{s}\nabla^{3}f(\hat{Y}_{r}^{y})(b_{\tau}(x),\sigma_{j,\tau}(x),\sigma_{j,\tau}(x))$
$\quad+\frac{1}{2}\sum_{k=1}^{m}\nabla^{4}f(\hat{Y}_{r}^{y})(\sigma_{j,\tau}(\mathscr{P}(x)),\sigma_{j,\tau}(\mathscr{P}(x)),\sigma_{k,\tau}(x),\sigma_{k,\tau}(\mathscr{P}(x)))\mathrm{d}r\Big]\mathrm{d}s\Big|$
$\leq\frac{1}{8}(2\|\nabla^{3}f\|\|\sigma_{\tau}(\mathscr{P}(x))\|_{HS}^{2}|b_{\tau}(\mathscr{P}(x))|+\|\nabla^{4}f\|\|\sigma_{\tau}(\mathscr{P}(x))\|_{HS}^{4})\tau^{2}.$
Similarly, the last four identities of the representation (2.13) imply that
$|\mathbb{E}\mathcal{R}_{3}(x,g_{\tau}(x))|$ $\leq\|\nabla^{2}f\||\mathscr{P}(x)-x||b_{\tau}(\mathscr{P}(x))|\tau,$
$|\mathbb{E}\mathcal{R}_{4}(x,g_{\tau}(x))|$ $\leq\|\nabla^{3}f\||\mathscr{P}(x)-x|\|\sigma_{\tau}(\mathscr{P}(x))\|_{HS}^{2}\tau,$
$|\mathbb{E}\mathcal{R}_{5}(x,g_{\tau}(x))|$ $\leq\|\nabla f\||b_{\tau}(\mathscr{P}(x))-b(x)|\tau,$
$|\mathbb{E}\mathcal{R}_{6}(x,g_{\tau}(x))|$ $\leq\|\nabla^{3}f\|\sum_{j=1}^{m}|\sigma_{j,\tau}(\mathscr{P}(x))-\sigma_{j}(x)|\tau.$
In combination the estimates in (3.14) with Assumption 3.2 and the estimate (3.3), we obtain
(3.19) $\Big|\sum_{i=1}^{6}\mathbb{E}^{\pi_{\tau}}[\mathbb{E}\mathcal{R}_{i}(x,g_{\tau}(x))|_{x=Y_{0}}]\Big|$ $\leq C\|\nabla^{4}f\|(1+\pi_{\tau}^{M}(|\cdot|^{\alpha_{1}}))\tau^{2},$
(3.20) $|\mathbb{E}^{\pi_{\tau}}\mathcal{A}_{\tau}f_{\varphi}(\mathscr{P}(Y_{0}))|$ $\leq C\|\nabla f\|\pi_{\tau}^{M}(|\cdot|^{\alpha_{2}})\tau^{2}.$
Then we conclude (3.15) by the above two estimates (3.19) and (3.20), together with the estimate (2.14).
Finally, we need to remark that $\pi^{M}_{\tau}(|\cdot|^{\max\{\alpha_{1},\alpha_{2}\}})<\infty$, as $\pi_{\tau}^{M}$ possesses all bounded $2p$-moments for any $p\in[1,p^{\star}]\cap\mathbb{N}_{+}$, which follows from the moments estimate of the MEM (3.10) in [12, Theorem 3.2].
(2) For the BEM (3.13), the associated intepolation process $(\hat{Y}_{t})_{t\in[0,\tau]}$ can be choosen as
(3.21) $\hat{Y}_{t}=\hat{Y}_{0}+\int_{0}^{t}b(Y_{0})\mathrm{d}s+\int_{0}^{t}\sigma(Y_{0})\mathrm{d}W(s),\quad t\in[0,\tau],$
i.e., $g_{\tau}(x):=x-b(x)\tau$ and $\tilde{g}_{\tau}(x)=x$, $x\in\mathbb{R}^{d}$. It is straightforward to verify that $\hat{Y}_{\tau}=Y_{1}-b(Y_{1})\tau=g_{\tau}(Y_{1})$, and that $\hat{Y}$ is continuous so that it is an Itô process.
Let $Y_{0}\sim\pi_{\tau}$. As $\pi_{\tau}$ is an invariant measure of (2.9), we have $Y_{1}\sim\pi_{\tau}$. Since $\hat{Y}_{0}=g_{\tau}(Y_{0})$ and $\hat{Y}_{\tau}=g_{\tau}(Y_{1})$, $\hat{Y}_{0}$ and $\hat{Y}_{\tau}$ share the same law. Then (2.11) implies $\mathbb{E}^{\pi_{\tau}}\mathcal{A}_{\tau}f_{\varphi}(\hat{Y}_{0})=0$.
Similarly to the above estimates of $\mathbb{E}\mathcal{R}_{i}(x,g_{\tau}(x))$ with $i=1,2,3,4$ for the MEM (3.10), we have
$|\mathbb{E}\mathcal{R}_{1}(x,g_{\tau}(x))|$ $\leq\frac{1}{4}(2\|\nabla^{2}f\||b(x)|^{2}+\|\nabla^{3}f\|\|\sigma(x)\|_{HS}^{2}|b(x)|)\tau^{2},$
$|\mathbb{E}\mathcal{R}_{2}(x,g_{\tau}(x))|$ $\leq\frac{1}{8}(2\|\nabla^{3}f\|\|\sigma(x)\|_{HS}^{2}|b(x)|+\|\nabla^{4}f\|\|\sigma(x)\|_{HS}^{4})\tau^{2}$
$|\mathbb{E}\mathcal{R}_{3}(x,g_{\tau}(x))|$ $\leq\|\nabla^{2}f\||b(x)|^{2}\tau^{2},$
$|\mathbb{E}\mathcal{R}_{4}(x,g_{\tau}(x))|$ $\leq\|\nabla^{3}f\||b(x)|\|\sigma(x)\|_{HS}^{2}\tau^{2}.$
Since $\tilde{g}_{\tau}(x)=x$, $x\in\mathbb{R}^{d}$, and there is no tame for the BEM (3.13), it is clear that
$|\mathbb{E}\mathcal{R}_{5}(x)|=|\mathbb{E}\mathcal{R}_{6}(x)|=0.$
Consequently,
$\Big|\sum_{i=1}^{6}\mathbb{E}\mathcal{R}_{i}(x)\Big|$ $\leq C(1+|b(x)|^{2}+|b(x)|\|\sigma(x)\|_{HS}^{2}+\|\sigma(x)\|_{HS}^{4})\tau^{2}$
(3.22) $\leq C(1+|x|^{2(\gamma+1)})\tau^{2}.$
In combination with the estimate (3.3) and the fact $\mathbb{E}^{\pi_{\tau}}\mathcal{A}_{\tau}f_{\varphi}(\hat{Y}_{0})=0$, we conclude (3.16).
Finally, we mention that $\pi^{B}_{\tau}(|\cdot|^{2(\gamma+1)})<\infty$, as $\pi_{\tau}^{B}$ possesses all bounded $2p$-moments for any $p\in[1,p^{\star}]\cap\mathbb{N}_{+}$, which follows from the moments estimate of the BEM (3.13) in [11, Lemma 10]. ∎
Remark 3.6.
Using the same arguments as in the proof of Theorem 3.5, we have the following sharp ergodic error estimate for the EM scheme with a unique invariant measure $\pi_{\tau}^{E}$, under the Lipschitz conditions as in [2, 4, 15]:
$\sup_{\|\varphi\|_{\mathcal{C}_{b}^{4}}\leq 1}|\pi_{\tau}^{E}(\varphi)-\pi(\varphi)|\leq C\tau.$
Remark 3.7.
If one can derive the regularity estimate (3.4) of the Stein equation solution for any $\varphi\in\mathcal{B}_{b}(\mathbb{R}^{d})$ or ${\rm Lip}(\mathbb{R}^{d})$, then the ergodic error estimates in Theorem (3.5) hold under total variation distance and Wasserstein distance, respectively.